OpenStudy (anonymous):

Use algebra to find the inverse of the function f(x) = 4x^11+6

OpenStudy (anonymous):

y = 4x^11+6 x = 4y^11 + 6 x-6 = 4y^11 x-6/4 = y^11 11th root of (x-6/4) = 11 root of y I am stuck on how to take the 11th root and solve for y

OpenStudy (anonymous):

That is the simplest form you can have.

OpenStudy (anonymous):

I thought you could simplify again to y...

OpenStudy (anonymous):

\[\sqrt[11]{\frac{ x-6 }{ 4 }}=y\]

OpenStudy (anonymous):

I am still getting it wrong

OpenStudy (anonymous):

OK, I got it correct. The problem was when I went to enter it, I left out () around the inside part